OpenStudy (anonymous):

4-log7(x+5)=3 solve for x

OpenStudy (anonymous):

-25/7 if you need an explanation i can do it

OpenStudy (anonymous):

i need an explanation please :D

OpenStudy (anonymous):

start by isolating the log. subtract 4 from each side to get -log(7(x+5)=-1. then subtract both sides to get log(7(x+5)=1. now distribute the 7 to get log(7x+35)=1. now convert it to exponential form to get 10=7x+35. can you get it from there?

OpenStudy (anonymous):

yup thanks :)

OpenStudy (anonymous):

np